General Announcement: - A. Galler informs on the LIMP-25 for next week: 1. Water: dirty traps maintenance ongoing in all tunnels but without interruptions, only changes in water pressure; 2. Limited AC and humidification control is expected; 3. Central hot water maintenance ongoing; 4. PLC updates ongoing in some areas (SA1/3 tunnels); 5. Last instrument racks ramping up next week; 6. No electricity interruptions are planned for next week. - M. Prollius informs that constructions in the campus / roads near the shaft buildings will be taking place in order to investigate water leakage. Please be careful when using the back entrances or biking in the campus. - O. Bilani reminds groups to inform her when companies are planned to carry work at XHEXP1. CTRL: - Nothing to report. DA: - Nothing to report. EEE: - Nothing to report. VAC: - Nothing to report. XPD: - Nothing to report. DET: - Nothing new from DET. ITDM: - Not present. XRO: - Not present. LAS: - All quiet at LAS. UNSYS: - Nothing to report. - MEA surveillance availability is limited for the next week, but shall not affect UNSYS. SPB: - Nothing to report. FXE: - Nothing to report. MID: - Nothing to report. HED: - Nothing to report. SCS: - Nothing to report. SQS: - Nothing to report. SXP: - Not present.